# WAKESET: A Large-Scale, High-Reynolds Number Flow Dataset for Machine Learning of Turbulent Wake Dynamics

### v1

@@ 5.3 Data Augmentation
To further increase the size and diversity of the dataset, data augmentation techniques were employed. Data augmentation is a common and effective strategy in machine learning to artificially expand the dataset by applying transformations to the existing data AugmentedDataRotation ; AugmentedDataGeneralisation . In the context of CFD simulations, augmentation can help improve the generalisation ability of ML models and reduce the risk of over-fitting to specific flow conditions or orientations. It is particularly valuable when generating the original data is computationally expensive, as is the case with high-fidelity CFD.
@@
    Rotation: The flow field data was rotated around the vertical axis (yaw) to generate new orientations while preserving the underlying flow physics. This is analogous to the rotation techniques used in image recognition AugmentedDataRotation , where rotating an image does not change its fundamental content but provides a different perspective. For each simulation with a non-zero turning angle, rotations were applied to create equivalent flow fields for the opposite turning direction. For example, a simulation with a +10-degree turning angle was rotated to create a corresponding -10-degree turning angle case.
@@
    Enhanced Robustness: The inclusion of rotated and flipped instances makes the models more robust to variations in the input data and reduces the likelihood of overfitting to specific orientations present in the original dataset AugmentedDataGeneralisation .
@@ 5.4 Data Availability
To maximise the impact of this research and contribute to the advancement of ML applications in fluid mechanics, the generated dataset has been made publicly available to the research community HuggingFace . The dataset, in the form detailed in Section 5.2 is hosted. This includes all data files (volumes and planes), loading functions for python use as well as PyTorch DataLoaders to integrated into ML workflows. The dataset is 480GB in size and can be accessed via a range of methods in the provided location HuggingFace .
@@ 6 Model Performance Benchmark
For the 2D task, three GAN architectures were selected to represent different subsets of the current GAN landscape. The first is the conditional deep convolutional GAN (cDCGAN) GAN_Flow1 ; GAN_Flow2 . The second is the Self-Attention GAN (SAGAN) zhang2019self and the third is the PatchGAN isola2017image . For the more computationally demanding task of 3D generation, the SAGAN, Wasserstein GAN with Gradient Penalty (WGAN-GP) WasserTraining ; WasserGAN and a custom 2D slice ingesting 2D3DGAN CFD_loss_3 were used.
@@ 6.2 Evaluation Metrics
The SSIM SSIM , from Eq. 3, is calculated between two windows of pixel values of the same size. In this formulation, $\mu_{x}$ denotes the pixel sample mean of $x$, $\mu_{y}$ the sample mean of $y$, $\sigma_{x}^{2}$ is the sample variance of $x$, $\sigma_{y}^{2}$ is the sample variance of $y$, $\sigma_{xy}$ is the sample covariance of $x$ and $y$. $c_{1}$ and $c_{2}$ are two variables introduced to stabilise the division with a weak denominator and are defined to be $c_{1}=(k_{1}L)^{2}$, $c_{2}=(k_{2}L)^{2}$ where $L$ is the dynamic range of the pixels and $k_{1}$ and $k_{2}$ are 0.01 and 0.03 respectively by default.
@@
The Fréchet Inception Distance (FID) FID_OG , provided in Eq. 4, is the standard metric used to assess the quality of images created by generative models, such as generative adversarial networks (GANs) and diffusion models. The FID compares the distribution of the generated images with the distribution of a set of real images (ground truths). A convolutional neural network, such as the inception architecture, produces high level features which are then compared. The Inception v3 model Szegedy_2016_CVPR was used in this work. The FID compares two sets of images, as represented by two multidimensional Gaussian distributions $\mathcal{N}(\mu,\Sigma)$ and $\mathcal{N}(\mu^{\prime},\Sigma^{\prime})$. The distance between the two distributions is calculated as the 2-Wasserstein distance between the two distributions. Higher distances (FID scores) between the distributions indicate a poorer generative model, whilst conversely, lower scores indicate better performance. A score of 0 would indicate a perfect model.
@@
The benchmark results for the two tasks are provided in Table 2. All architectures were trained for a period of 200 epochs, with a batch size of 16 instances (2D or 3D). The training was undertaken in PyTorch on a dual NVIDIA RTX A6000 system, equipped with 256GB of RAM. In training, an epoch-batch loop was implemented. The Adam optimiser was used, and the GAN Binary Cross Entropy (BCE) loss was augmented with a structural flow metric. This was the Gradient Mean Squared Error (GMSE) loss cooper2024generalised . This loss is specifically designed to operate on computational fluid dynamics data, such as the WAKESET. In addition to the metrics above, are the inference time $t$ and the model size upon inference $M_{s}$ to give an indication of the computational efficiency.
@@
2D $512^{2}$] cDCGAN GAN_Flow1 47.24 0.998 43.01 0.015 4.51 0.422
SAGAN zhang2019self 18.70 0.536 329.51 0.283 15.80 1.203
PatchGAN isola2017image 46.26 0.997 54.43 0.018 4.51 0.423
3D $128^{3}$] SAGAN zhang2019self 29.06 0.602 27.99 0.081 95.97 10.39
WGAN-GP WasserGAN 15.62 0.478 341.98 0.312 14.08 1.21
2D3DGAN CFD_loss_3 31.62 0.752 285.53 0.113 1.48 0.306
@@ 7 Discussion
The development of a novel, large-scale, high-fidelity CFD dataset is specifically designed to address the critical need for comprehensive training data in machine learning applications for fluid mechanics that are highly turbulent, sufficiently large and focused on a practical engineering problem. The dataset was developed based on the complex hydrodynamics of underwater vehicle interactions, using the scenario of an AUV recovering within an XLUUV. The methodology began with a detailed hydrodynamic analysis of this specific scenario IEEE , then before it was systematically generalised and expanded from this initial investigation to create a versatile resource for training a variety of ML models (Section 4.3) by ensuring relevant physical effects and flow structures were captured. The dataset development process demonstrated:
It is observed that the new dataset offers the second highest number of training instances Huang , the highest upper Reynolds number Ahmed_2016 , and comprehensive 3D data, features that significantly enhance its applicability for training advanced ML architectures in fluid mechanics. The large number of instances (4,364 after augmentation) allows for the training of more complex, data-intensive ML models (e.g., deep neural networks with high parameter counts) whilst mitigating the risk of overfitting and potentially promoting better generalisation to unseen flow conditions. The high Reynolds limit (up to $1.09\times 10^{8}$) ensures that the dataset represents highly turbulent flows, characteristic of many full-scale engineering applications. This enables ML models to learn features relevant to these challenging, practical regimes, which are often underrepresented in existing datasets focused on lower Re or canonical flows. The 3D nature of the data is crucial for capturing the inherently three-dimensional physics of most real-world fluid dynamic problems. Including complex vortical structures, flow separation, and anisotropic turbulence. This allows ML models to learn and predict these complex spatial features, which are often lost or oversimplified in 2D datasets MegaFlow2D ; McConkey ; AirfRANS ; Huang . Collectively, these attributes position the dataset as a valuable resource for developing and benchmarking ML models capable of handling the complexities of practical engineering flows.

### v2

@@ 5.3 Data Augmentation
To further increase the size and diversity of the dataset, data augmentation techniques were employed. Data augmentation is a common and effective strategy in machine learning to artificially expand the dataset by applying transformations to the existing data [35, 33]. In the context of CFD simulations, augmentation can help improve the generalisation ability of ML models and reduce the risk of over-fitting to specific flow conditions or orientations. It is particularly valuable when generating the original data is computationally expensive, as is the case with high-fidelity CFD.
@@
    Rotation: The flow field data was rotated around the vertical axis (yaw) to generate new orientations while preserving the underlying flow physics. This is analogous to the rotation techniques used in image recognition [35], where rotating an image does not change its fundamental content but provides a different perspective. For each simulation with a non-zero turning angle, rotations were applied to create equivalent flow fields for the opposite turning direction. For example, a simulation with a +10-degree turning angle was rotated to create a corresponding -10-degree turning angle case.
@@
    Enhanced Robustness: The inclusion of rotated and flipped instances makes the models more robust to variations in the input data and reduces the likelihood of overfitting to specific orientations present in the original dataset [33].
@@ 5.4 Data Availability
To maximise the impact of this research and contribute to the advancement of ML applications in fluid mechanics, the generated dataset has been made publicly available to the research community [21]. The dataset, in the form detailed in Section 5.2 is hosted. This includes all data files (volumes and planes), loading functions for python use as well as PyTorch DataLoaders to integrated into ML workflows. The dataset is 480GB in size and can be accessed via a range of methods in the provided location [21].
@@ 6 Model Performance Benchmark
For the 2D task, three GAN architectures were selected to represent different subsets of the current GAN landscape. The first is the conditional deep convolutional GAN (cDCGAN) [31, 45]. The second is the Self-Attention GAN (SAGAN) [46] and the third is the PatchGAN [30]. For the more computationally demanding task of 3D generation, the SAGAN, Wasserstein GAN with Gradient Penalty (WGAN-GP) [25, 4] and a custom 2D slice ingesting 2D3DGAN [44] were used.
@@ 6.2 Evaluation Metrics
The SSIM [42], from Eq. 3, is calculated between two windows of pixel values of the same size. In this formulation, $\mu_{x}$ denotes the pixel sample mean of $x$, $\mu_{y}$ the sample mean of $y$, $\sigma_{x}^{2}$ is the sample variance of $x$, $\sigma_{y}^{2}$ is the sample variance of $y$, $\sigma_{xy}$ is the sample covariance of $x$ and $y$. $c_{1}$ and $c_{2}$ are two variables introduced to stabilise the division with a weak denominator and are defined to be $c_{1}=(k_{1}L)^{2}$, $c_{2}=(k_{2}L)^{2}$ where $L$ is the dynamic range of the pixels and $k_{1}$ and $k_{2}$ are 0.01 and 0.03 respectively by default.
@@
The Fréchet Inception Distance (FID) [27], provided in Eq. 4, is the standard metric used to assess the quality of images created by generative models, such as generative adversarial networks (GANs) and diffusion models. The FID compares the distribution of the generated images with the distribution of a set of real images (ground truths). A convolutional neural network, such as the inception architecture, produces high level features which are then compared. The Inception v3 model [39] was used in this work. The FID compares two sets of images, as represented by two multidimensional Gaussian distributions $\mathcal{N}(\mu,\Sigma)$ and $\mathcal{N}(\mu^{\prime},\Sigma^{\prime})$. The distance between the two distributions is calculated as the 2-Wasserstein distance between the two distributions. Higher distances (FID scores) between the distributions indicate a poorer generative model, whilst conversely, lower scores indicate better performance. A score of 0 would indicate a perfect model.
@@
The benchmark results for the two tasks are provided in Table 2. All architectures were trained for a period of 200 epochs, with a batch size of 16 instances (2D or 3D). The training was undertaken in PyTorch on a dual NVIDIA RTX A6000 system, equipped with 256GB of RAM. In training, an epoch-batch loop was implemented. The Adam optimiser was used, and the GAN Binary Cross Entropy (BCE) loss was augmented with a structural flow metric. This was the Gradient Mean Squared Error (GMSE) loss [20]. This loss is specifically designed to operate on computational fluid dynamics data, such as the WAKESET. In addition to the metrics above, are the inference time $t$ and the model size upon inference $M_{s}$ to give an indication of the computational efficiency.
@@
2D $512^{2}$] cDCGAN [31] 47.24 0.998 43.01 0.015 4.51 0.422
SAGAN [46] 18.70 0.536 329.51 0.283 15.80 1.203
PatchGAN [30] 46.26 0.997 54.43 0.018 4.51 0.423
3D $128^{3}$] SAGAN [46] 29.06 0.602 27.99 0.081 95.97 10.39
WGAN-GP [4] 15.62 0.478 341.98 0.312 14.08 1.21
2D3DGAN [44] 31.62 0.752 285.53 0.113 1.48 0.306
@@ 7 Discussion
The development of a novel, large-scale, high-fidelity CFD dataset is specifically designed to address the critical need for comprehensive training data in machine learning applications for fluid mechanics that are highly turbulent, sufficiently large and focused on a practical engineering problem. The dataset was developed based on the complex hydrodynamics of underwater vehicle interactions, using the scenario of an AUV recovering within an XLUUV. The methodology began with a detailed hydrodynamic analysis of this specific scenario [17], then before it was systematically generalised and expanded from this initial investigation to create a versatile resource for training a variety of ML models (Section 4.3) by ensuring relevant physical effects and flow structures were captured. The dataset development process demonstrated:
It is observed that the new dataset offers the second highest number of training instances [28], the highest upper Reynolds number [2], and comprehensive 3D data, features that significantly enhance its applicability for training advanced ML architectures in fluid mechanics. The large number of instances (4,364 after augmentation) allows for the training of more complex, data-intensive ML models (e.g., deep neural networks with high parameter counts) whilst mitigating the risk of overfitting and potentially promoting better generalisation to unseen flow conditions. The high Reynolds limit (up to $1.09\times 10^{8}$) ensures that the dataset represents highly turbulent flows, characteristic of many full-scale engineering applications. This enables ML models to learn features relevant to these challenging, practical regimes, which are often underrepresented in existing datasets focused on lower Re or canonical flows. The 3D nature of the data is crucial for capturing the inherently three-dimensional physics of most real-world fluid dynamic problems. Including complex vortical structures, flow separation, and anisotropic turbulence. This allows ML models to learn and predict these complex spatial features, which are often lost or oversimplified in 2D datasets [43, 36, 10, 28]. Collectively, these attributes position the dataset as a valuable resource for developing and benchmarking ML models capable of handling the complexities of practical engineering flows.